erhält Unterricht in/von

\underline{Fach ID} \underline{Klassen ID}

Personalnummer

Deutsch  aW_24-28 0001
Deutsch  bW_24-28 0001
Deutsch  cW_24-28 0001
Deutsch dP_24-28 0001
Deutsch  eW_24-28 0001
Deutsch fP_24-28 0001
Französisch aW_24-28 0002
...  ...  ...